A ball is dropped from top of the building.The ball takes 0.5s to fall past 3m of a window some distance fro the top of the building. If the speedof the ball at top and bottom of the vindow are’ U’ and’V’ respectively then ‘U-V’=?, V+V=? UV=? 

Grade:12th pass

4 Answers

Pranav Dabhade
7 years ago
V2-U2=2g(3) also, V=U+g(0.5) i.e V-U=0.5g
(V+U)(V-U)=6g
(V+U)(0.5g)=6g
V+U=12,U-V=-4.9, Q. Must be modified as,
A ball is dropped from top of the building.The ball takes 0.5s to fall past 3m of a window some distance fro the top of the building. If the speedof the ball at top and bottom of the vindow are’ U’ and’V’ respectively then ‘U-V’=?, V+U=? UV=?
Aman
7 years ago
Let initial velocity at top of window be 'u' and velocity at bottom be ' v' 
Use S = ut + 1/2at2   where u is initial velocity a =g(10) t =0.5  S =3 ,. We get value of u as 3.5m/S. Now use v = u + at  where u is initial velocity and a =g,t =0.5 you will get the value of v now use it to find u-v ,v+v ,u.v
Shivansh
4 years ago
Here V, S, and U is given... 
So,  use 
              V=U+GT     ;         (U+V) t=2S
              V-U=GT       ;
NOW PUT THE VALUES YOU WILL GET PERFECT ANSWERS.... U+V=12 AND U-V=-4.9    
Yash Chourasiya
3 years ago
Hello Student

Let initial velocity at top of window be 'U' and velocity at bottom be ' V'
V2 – U2 = 2aS
V2- U2= 2g(3) ….........(1)
also,V = U + at
V = U + g(0.5)
V – U = 0.5g........(2)
Equation (1) can also be written as
(V+U)(V-U)=6g
Putting value of V – U from eq. (2), we get
(V+U)(0.5g)=6g
V + U = 12 m/sec …...(3)
In equation (2) putting value of g, and we get
U – V = -4.9 m/sec …....(4)
From equation (3) & (4), we get
UV = 29.9975 = 30 aprox

I hope this answer will help you.